If a plane flies eastward for four hours with a 60 mph tailwind and returns in five hours, what was the total distance traveled?
(Hint: the rate of travel eastbound is S + 60 and westbound is S - 60. S is the speed of the plane in still air.)


Answer by rapaljer(4671)

D = R*T
Let S = Rate in still air
S + 60 = Rate eastbound
S - 60 = Rate westbound
4 hours = Time eastbound
5 hours = Time westbound

Distance traveled eastbound = Distance traveled westbound.
4(S + 60) = 5 (S - 60)
4S + 240 = 5S - 300

Add -4S to each side to get all variables to the right side
4S - 4S + 240 = 5S - 4S - 300
240 = S - 300

Add 300 to each side:
240 + 300 = S - 300 + 300
540 MPH = S